OpenStudy (anonymous):

evaluate limit as x approaches zero from the right. sin (x) ln (3x).

OpenStudy (anonymous):

Have you learned the Squeeze/Sandwhich theorem?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

It's not the whole probelm that i need help with it's just the part at the very end where you have sin (x)/x = 1

OpenStudy (anonymous):

Wait so is your question with sin(x) ln(3x) or sin(x) / x?

OpenStudy (anonymous):

it's dealing with that problem, just the very end of it. i 'm just wondering how the limit as x approaches 0 from the right of sin(x)/x is 1

OpenStudy (anonymous):

Its a Squeeze Theorem proof and its a geometrical one: http://www.youtube.com/watch?v=Ve99biD1KtA

OpenStudy (anonymous):

to understand that you need to know l'hopitals. Thats why in calc 1 they call these "Special limits"

OpenStudy (anonymous):

ok i get you so far...

OpenStudy (anonymous):

if you have the form \[\frac{0}{0}\], you can find the limit by take the derivative of the top and the derivative of the bottom and evaluate the limit again until you do not get a indeterminate form

OpenStudy (anonymous):

so taking the derivative you get \[\frac {cos(x)}{1}\] as x goes to 0, this equals 1

OpenStudy (anonymous):

that is why sinx/x equals 1

OpenStudy (anonymous):

Oooo! ok thanks so much! One more question. How does the limit as x approaches 0 of cot (2x) equal infinity?

OpenStudy (anonymous):

cot(2x) is 1/tan(2x) as x->0, tan(2x) is going to tan(0) which is 0. so you have 1/0

OpenStudy (anonymous):

this you cannot use the l'hopitals quite yet one way i cant think of is using double angle stuff

OpenStudy (anonymous):

1/0 is undefined not infinity

OpenStudy (anonymous):

\[\frac{cos(2x)}{sin(2x)}\]

OpenStudy (anonymous):

That's what I thought too but im doing homework on this online thing and they said that 0 was wrong and it was infinity so I j

OpenStudy (anonymous):

anted to know why

OpenStudy (anonymous):

no what i'm saying is that you cannot assume that it is infinity by seeing that it's undefined... there is a chance it is infinity though

OpenStudy (anonymous):

okey doke

OpenStudy (anonymous):

in this case it is infinity